OpenStudy (anonymous):

Guys I Seriously Don't Get This Please Help!! -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

Meadling too!!

OpenStudy (anonymous):

Easy!! multiply the top eaquasion by -3, so you get 9x-21y=48, and -9x+5y=16. then cancel out the X to get -16y=64 then solve for Y

OpenStudy (anonymous):

how would you create a equivalent equation?

OpenStudy (anonymous):

Do you know this subject material/ Do you know what an Equivalent Equation is??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sortof

OpenStudy (anonymous):

So then tell me what you think you need to do to find the Equivalent Equations

OpenStudy (anonymous):

umm write an equation that is equal to the first equation ? idk

OpenStudy (anonymous):

you do the same thing to find X and then you plug in the number to find the solution. The solutions should be Equivalent. Thats your answer. Tell me what you get for X and we will go from there

OpenStudy (anonymous):

the first one?

OpenStudy (anonymous):

wait what?? Can you just tell me the answer lol

OpenStudy (anonymous):

multiply the top equation by -5 and the bottom by 7 to cancel out the Y. Tell me what you get for the new equations

OpenStudy (anonymous):

this is prolly so wrong but i got 15x=80 -63x=112

OpenStudy (anonymous):

then combine

OpenStudy (anonymous):

tell me whatcha get

OpenStudy (anonymous):

are those right?

OpenStudy (anonymous):

wait but how do you combine them

OpenStudy (anonymous):

you do 15 -63 to get -48X=192, then X=-4

OpenStudy (anonymous):

and what did you get for Y=_??

OpenStudy (anonymous):

what does x equal again?

OpenStudy (anonymous):

wait im totally confused @Da_Homie

OpenStudy (anonymous):

X= -4

OpenStudy (anonymous):

do the same thing to get Y

OpenStudy (anonymous):

or just plug in X and solve for Y

OpenStudy (anonymous):

once you get Y then we can solve the problem

OpenStudy (anonymous):

Tell me whatcha get fir Y

OpenStudy (anonymous):

x=-4 for both equations?

OpenStudy (anonymous):

yep, plug it in and solve for Y

OpenStudy (anonymous):

Plug it into both equations so we can see if you got X=-4 correct

OpenStudy (anonymous):

for the first one i got y=-2

OpenStudy (anonymous):

ok, now do the same for the next, and if you get Y=-2 also then X is correct

OpenStudy (anonymous):

i got y=-4

OpenStudy (anonymous):

so your X is wrong then... Your gonna need to redo X

OpenStudy (anonymous):

how do we do that

OpenStudy (anonymous):

multiply the top equation by 5

OpenStudy (anonymous):

multiply the bottom equation by -7

OpenStudy (anonymous):

what would that be then?

OpenStudy (anonymous):

-15X+35Y=80

OpenStudy (anonymous):

49X-35Y=-112

OpenStudy (anonymous):

Ys now cancel out

OpenStudy (anonymous):

so -15x=80 and 49x=-112

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

soo what now

OpenStudy (anonymous):

darn...the X comes out to a decimal. Lets start with Y lol

OpenStudy (anonymous):

jus hold on ill figure it out for u

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Your sure these are the right numbers????

OpenStudy (anonymous):

yes i copy and pasted it lol

OpenStudy (anonymous):

well... both of our numbers are gonna be decimals :(... that means our job is gonna get confusing

OpenStudy (anonymous):

oh :(

OpenStudy (anonymous):

unless were doing it wrong...?

OpenStudy (anonymous):

i dunno... everything comes to a decimal :(

OpenStudy (anonymous):

one of my friends might know @SolomonZelman

OpenStudy (anonymous):

you would have to wait for him to get online tho

OpenStudy (anonymous):

ok thats fine

OpenStudy (anonymous):

srry bud

OpenStudy (anonymous):

its ok i hate this test

OpenStudy (anonymous):

lol i getcha, same

OpenStudy (anonymous):

could you help me with another problem tho?

OpenStudy (anonymous):

x + 2y < -5 x - y > 0 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (anonymous):

on a number line right?

OpenStudy (anonymous):

no on a graph ...

OpenStudy (anonymous):

oh yes, i think i remember these, with the shading right??

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

have you done these before

OpenStudy (anonymous):

no ..

OpenStudy (anonymous):

darn

OpenStudy (anonymous):

i havent in a long time

OpenStudy (anonymous):

:// crapp

OpenStudy (anonymous):

we might be able to figure it out tho, do you know somewhat what you are doing??

OpenStudy (anonymous):

yeah i think you have to solve for y first

OpenStudy (anonymous):

then figure out the solutions

OpenStudy (anonymous):

wait no scratch the first part your supposed to find the solutions first

OpenStudy (anonymous):

do you know how to go about doing that???

OpenStudy (anonymous):

i think its the same way you were trying to to do with this problem ^

OpenStudy (anonymous):

ok so lets do it

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

do multiplly the bottom by 2 and cancel out the Y

OpenStudy (anonymous):

no you solve for y first

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

then solve for Y lol

OpenStudy (anonymous):

thats what im doing lol

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

wait lol do you subtract x from both sides ?

OpenStudy (anonymous):

multiply the bottom by -1, then the Xs will cancel out

OpenStudy (anonymous):

what do you mean multiply the bottom??

OpenStudy (anonymous):

the bottom equation needs to be multiplied by -1 so the Xs can cancel out

OpenStudy (anonymous):

multiply what?

OpenStudy (anonymous):

THE BOTTOM EQUATION lol

OpenStudy (anonymous):

NO multiply WHAT on the bottom equation?

OpenStudy (anonymous):

the whole equation X and Y are multiplied by -1

OpenStudy (anonymous):

like this -1( -9x+5y=16)

OpenStudy (anonymous):

BOTTOM

OpenStudy (anonymous):

that is the top

OpenStudy (anonymous):

that is the bottom one

OpenStudy (anonymous):

but it really doesnt matter, so that'l work ;)

OpenStudy (anonymous):

but that just increases everything